40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 202 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  7826   Fri Dec 14 01:42:53 2012 DenUpdateSUSTT angle of incidence

I've estimated max possible angle of incidence on TT if we allow 20mm tolerance for the beam size and 5 mm tolerance for spot location on the mirror. It turns out to be

alpha = 43 degrees

So we need to cut the central part of the bracket. Then the max possible angle of incidence will be

alpha = 63 degrees

 

DSC_4791.jpeg

We can start the vent on Monday and use TT with an old bracket for yaw damping and later during the week we can install the brackets after they will be baked.

  7825   Thu Dec 13 21:21:34 2012 DenUpdatePEMseismometers

Quote:

 

 Looking at the PEM BLRMS, I noticed that the GUR1Z channel had a much reduced microseism compared to the GUR1X. Looking at the BLRMS screens everything seems ON, although its a mess (too many filters in the banks, etc. - clean this up, PEM people).

 Looking at the Z channel in DTT, I see that the Z spectra looks double high pass filtered below ~1 Hz.     Needs some attention in the daytime.

From Den and Ayaka's elog entry from Nov 29, its clear that this problem is there at that time. It seems that the seismometer was not even hooked up before then. Perhaps Tara returned the seismometer around Thanksgiving and then someone here hooked it up but neglected to log this work? If so, please make an elog now describing the installation of this sensor at the 40m and log any future work which takes place at the 40m lab even if you think it is inconsequential.

 Yesterday I wanted to recenter Guralps. I turned them off, understood that would be able to center them because we do not have power cable to Guralp box from Tara yet and turned them back on.

I've switched Guralp cables and spectrums are fine now.

Attachment 1: gur.pdf
gur.pdf
Attachment 2: gur_fix.pdf
gur_fix.pdf
  7824   Thu Dec 13 18:06:59 2012 AyakaUpdateWienerFilteringNew microphone mounts

 Yesterday, I made new mounts for microphones.

IMG_0129.JPGIMG_0130.JPG

I glued a microphone on a pedestal. The cables are attached loosely so that its tension does not make any noise.
At the bottom of the mount, I attached the surgical tube forming a ring by double-side tape so that it damps the seismic vibration.

I made 6 mounts and these are all on the AS table now.
I took some data of XARM signal controlled by AS.

My plan is to find/set an upper limit on acoustic coupling noise in AS signal.
The acoustic noise can be estimated by the Wiener filter, but it is not accurate because it may see residual correlation between AS and microphone signals that should be 0 when the data is long enough.
I will find/set an upper limit by the analysis based on Neuman-Pearson criterion, that is analog of a stochastic GW background search.
If I can find the acoustic coupling noise should be below the shot noise, I am happy. If not, some improvements may be needed someday.

  7823   Thu Dec 13 17:24:53 2012 AyakaUpdatePSLPMC calibration for MC_F calibration

 I calibrated MC_F signal into Hz/rtHz unit using the transfer function from MC_F to PMC feedback signal.

Here is the diagram:
  MC-PMCservo.jpg

n_mcf is MC_F signal we can get at dtt. I measured n_pmc/n'_mcf using SR.

TF_mcf-pmc.png

Other information I used:

G_out = 2.49/123.49 (see the document D980352-E01-C)

Fout has 1 pole at 10 Hz (see the document D980352-E01-C)

A_pzt = 371e+6/63 [Hz/V] (see elog)

F_wt has 1 pole at 100 Hz and 1 zero at 10 Hz.

Then, calibration transfer function of H is fitted as 1e+9/f [Hz/V]:

H.png

Then, the calibrated spectrum of MC_F is below:

 MCF_calib.pdf

This calibration have about 20 % error.
Compared to the spectrum in Jenne's paper (elog), above 20 Hz it seems to be laser frequency noise. But now we have extra unknown noise below 10 Hz.

Note: calibration value of laser's PZT is ~ 1MHz/V. This is reasonable compared to the data sheet of the laser. (This is calculated by combining result of H and transfer function of the circuit box1 and FSS.)

 laser.png

Attachment 6: calib.zip
  7822   Thu Dec 13 04:42:32 2012 DenUpdateSUSITMX local damping

Tonight we've noticed that ITMX local damping was kicking the optics. This happened because LR shadow sensor was not working. In ~30 minutes it started to work again. Evan and I were working on installation, moving and focusing cameras and locking prcl and mich. We've installed a camera on BSC and plugged it in to PSL_SPARE input.

I'm not sure that this can be correlated to ITMX LR shadow sensor behaviour.

 

lrsen.png

  7821   Thu Dec 13 04:29:34 2012 DenUpdateSUSTT angle of incidence

I think the angle of incidence on TT inside BSC will be too large because of eddy current damping brackets. I've measured max possible angle of incidence

  Max angle of incidence, degrees
No bracket 72
Original bracket 45
New bracket (with no screws for tiny yaw magnets) 52

This means that we do not have too much range and there is a probability that 45 degree incident beam will start clipping. I think we should just cut off the central part of the bracket. We do not need it anyway, our eddy current damping due to corner magnets is good enough.

I've left the brackets near the laptop in the clean room.

  7820   Thu Dec 13 03:20:48 2012 DenUpdateLSCbeam inside DRMI is clipping on PR3 Tip-Tilt

Quote:

 

 I've made snapshots of PR2, PRM, ITMY and ITMX mirrors. Power buildup recycling gain (POWER BS / POWER PRM) was equal to 3-4.

           

 We've looked at PR2 face camera when PRM, BS and one of the ITMs were aligned. We saw an extra beam at PR2 when ITMX was aligned (right plot). This spot stays on the PR2 when prcl is locked.

PR2_ITMX.png   PR2_ITMY.png

Then we looked at PR3 transmission mirror and saw that the main beam is not on the edge of the mirror. Secondary beam is clipping on the mirror mount of PR3 that we see on BS_PRM camera.

PR3_LOCK.png

Measured beam spot positions:

Optics Pitch, mm Yaw, mm
ITMX 5.6 1.5
ETMX -1.5 1.5
ITMY 4.8 -1.5
ETMY -1.4 5.6
PRM 2.7 4.1

"+" for pitch means that the beam is too high, "-" too low

"+" for yaw means that the beam is left if you look from the back, "-" is right

Beam spots were measured using x, y arm and prcl locking to the carrier.

  7819   Thu Dec 13 01:49:53 2012 ranaUpdatePEMseismometers

Quote:

I've installed Guralp readout box back and it turned out that it does not work with voltage provided from the rack (+13.76 0 -14.94).  +/-12 voltage regulators inside the box convert it to -0.9 0 -12. I've connected the box to +/-15 DC voltage supply to measure seismic motion at the ETMY table. Readout box works fine with +/- 15.

Seismic noise on the ETMY table measured to be a few times higher then on the floor in horizontal direction in the frequency range 50 - 200 Hz. Attached are compared spectrums of X, Y and Z motions.

 Looking at the PEM BLRMS, I noticed that the GUR1Z channel had a much reduced microseism compared to the GUR1X. Looking at the BLRMS screens everything seems ON, although its a mess (too many filters in the banks, etc. - clean this up, PEM people).

 Looking at the Z channel in DTT, I see that the Z spectra looks double high pass filtered below ~1 Hz.     Needs some attention in the daytime.

From Den and Ayaka's elog entry from Nov 29, its clear that this problem is there at that time. It seems that the seismometer was not even hooked up before then. Perhaps Tara returned the seismometer around Thanksgiving and then someone here hooked it up but neglected to log this work? If so, please make an elog now describing the installation of this sensor at the 40m and log any future work which takes place at the 40m lab even if you think it is inconsequential.

Attachment 1: a.pdf
a.pdf
  7818   Wed Dec 12 20:22:03 2012 JamieUpdateComputer Scripts / Programsilluminators fixed and added to VIDEO screen

I fixed the illuminator setup.  ETMY was not hooked up, and the screen wasn't configured quite right.  The ITMX illuminator still needs to be hooked up to the vertex switch.

I made an updated illuminator script that works more like the videoswitch scripts, with a saner interface, and is located here:

/opt/rtcds/caltech/c1/scripts/general/illuminator

I also fixed up the illuminator MEDM interface a bit and added it to the VIDEO screen:

video.png

While I was at it, I cleaned up the sitemap a bit:

sitemap.png

I hope everyone won't be too confused.

  7817   Wed Dec 12 17:26:47 2012 RijuUpdate Testing AG4395A+GPIB

I repeated my experiment to get noise level. To get that I disconnected the bandpass filter SBP-10.7  from channel A of network analyzer AG4395A and terminated both the open ends (open end of filter and open end of channel A) with 50ohm terminator.

Reference level had been corrected, signal and noise data had been collected separately w.r.t that level.

Command for GPIB:   ./netgpibdata.py -i 192.168.113.105 -d AG4395A -a 10 -f filename

The result is as follows

 

Attachment 1: TFbandpassfilter.pdf
TFbandpassfilter.pdf
  7816   Wed Dec 12 16:52:12 2012 JenneUpdateAlignmentPR2_face, PR3_back cameras in place

I have setup cameras looking at the back of PR3 (through the north viewport on the MC chamber) and the face of PR2 (through the north viewport on the ITMX chamber). We would like a view of the face of PR3, but that isn't possible without placing another in-vac mirror.  The best we can do is the current PRM_BS camera setup, which sees a small portion of the PR3 face.  Most of the face is obscured by the PRM itself. 

I have taken images with the PRM misaligned.  The spot near the top of PR2 is the first reflection from the pitch-misaligned PRM, so it should be ignored for the purposes of trying to see the straight-shot, no PRM beam.

Images are taken with my videocapture50 script, in ..../scripts/general/videoscripts.  This takes 10 sets of 50 images and saves them.  Then ImageBkgndSubtractor.m located in the same folder takes the images (you must edit the beginning of the script to tell it where the images are), averages the noBeam images (PSL shutter closed), and averages the withBeam images, and subtracts them.  Results below:

PR2_face_12Dec2012_SpotImage.pngPR2_face_12Dec2012_SpotImage_pixelsTimes3.png

PR3_back_12Dec2012_SpotImage.pngPR3_back_12Dec2012_SpotImage_pixelsTimes3.png

  7815   Wed Dec 12 14:59:49 2012 ManasaConfiguration40m UpgradingNew tip-tilts layout in BSC

I have updated the BSC layout to include the new tip-tilts. The bigger footprints of the tiptilts are on in the way of the existing PRM oplev path. So I have recalculated new PRM oplev paths. The proposed layout requires a new oplev mirror to be included.

 

bs.png

  7814   Wed Dec 12 11:49:05 2012 AyakaUpdateLSCAcoustic noise in POX and AS error signal

[Koji, Ayaka]

Last night, I injected acoustic noise at POX table and AS table with oplev controls on (LPF is on).

 

1. acoustic noise at the POX table

I set the microphones and speakers at the POX table and see the acoustic coupling.

POX_noise4.pdf

I could see slight change around 40 Hz. This can be caused by the oplev feedback loop because the speaker was on the same table as the ITMX oplev.

 

2. acoustic noise at the AS table

I controlled XARM with AS error signal and set the microphones and speaker on the AS table.

AScontXARM_acoustic.pdf

The resonance a 200 Hz seemed to be enhanced. But still we are not sure that it is caused by acoustic noise. Because this resonance is enhanced when the OL gain is high, and the gain adjustment was so critical that this resonance was easily enhanced even when the acoustic noise is not injected. And sometimes it has gone away.

  7813   Wed Dec 12 11:04:45 2012 ManasaConfiguration40m UpgradingEndtable upgrade for auxiliary green laser : ETMX layout on new table

I have updated the layout to fix all the issues brought up. The last couple of 2" green steering mirrors will hold the PZTs for input steering. I will update with the list of optical components that we will be ordering for this layout. The ETMY endtable layout will be similar to this one, except that we will have IPANG setup at the empty space in the right top corner.

ETMX_endtable_New_Model.png

  7812   Tue Dec 11 21:53:37 2012 AyakaUpdateLSCOplev update for improving sensitivity

[Rana, Ayaka]

The BS oplev pitch feedback came back.

OL_BSpit.pdf

The problem was that 300^2:0 filter was off. And I turned on all the low pass filters (ELP35), then the oplev servo does not seem to inject big noise into the arms as long as I see the spectra of POY and POX. These low-pass filters will be modified tomorrow so that the acoustic coupling noise is minimized.

BSoplevservo.png

  7811   Tue Dec 11 19:51:36 2012 KojiUpdatePSLPMC gain was too low / EPICS alerting value for PMC updated

[Ayaka, Koji]

Ayaka pointed out that the PMC has too low unity gain frequency. We checked the history of "C1:PSL-PMC_GAIN"
and found that the gain was minimum from the Friday night. It was returned to nominal gain of 10.

The PMC screen had the gain status indicator always red. This was because C1:PSL-STAT_PMC_NOM_GAIN was 2 instead of 10.
This was fixed by the following command.

ezcawrite C1:PSL-STAT_PMC_NOM_GAIN 10

This will be recorded by the snapshot in an hour.

Another annoying false alerm on the PMC screen was the PMC transmission monitor.
In order to fix this, the following commands were executed.

ezcawrite C1:PSL-PMC_PMCTRANSPD.LOLO 0.75
ezcawrite C1:PSL-PMC_PMCTRANSPD.LOW 0.8
ezcawrite C1:PSL-PMC_PMCTRANSPD.HIGH 0.9
ezcawrite C1:PSL-PMC_PMCTRANSPD.HIHI 0.95

Also the corresponding EPICS database (/cvs/cds/caltech/target/c1psl/psl.db) has been updated accordingly.

  7810   Tue Dec 11 11:40:07 2012 Manasa, AyakaUpdatePSLPMC drift

[Manasa, Ayaka]

I found that MC got unstable this morning. This is caused by the drift of PMC. The transmission of PMC was going down and eventually unlocked PMC.

PMCdrift121211.pdf

We adjusted 'Slow Actuator Adjust' in FSS and now the PMC is locked with transmission of ~ 0.735.
Also we aligned the MC to be locked. Now it is locked with transmission of ~ 0.5 with WFS and MCL on.

  7809   Tue Dec 11 10:09:04 2012 AyakaHowToPEMcables needs care

Quote:

How NOT to:

The janitor can not clean in areas like this. He may only steps on these cables accidentally as he dust wiping our chambers.

 Sorry for the mess. I fixed it.

  7808   Tue Dec 11 09:31:47 2012 AyakaUpdateLSCOplev update for improving sensitivity

 Motivation

We observed that oplev servos affect the arm spectra badly (elog #7798). Some of them are fixed, but still they inject noise into the arms.
So I tried to turn the oplevs off and to see the acoustic noise effect. However, the mirrors moves so much that the signal does not seem to be linear any more, and the noise spectrum of arms changes especially around 60 - 100 Hz as you can see the spectrogram of YARM error signal below. This makes it difficult to find acoustic coupling noise. Therefore, I tried to fix the oplev servos so that the noise spectra do not get worse when the oplev servos are on.
POYspectrogram_nonoise.png

Checking oplev UGFs

I checked the oplev open loop transfer functions. The UGFs of oplevs are all around 1-3Hz and phase margin looks enough except the BS oplev.
The gain of the BS oplev OLTF has so low that the signal is not fed back. Moreover, there is much phase delay in the BS feedback loop than the others'.
The counts of BS oplev sum is not changed so much for this 4 months, so the oplev beam seems to hit the BS correctly.
I am not sure what makes difference.
ETMYoplev.pdfOL_BSoplevPIT.pdf

 BSoplev.pdf

 

Clipped oplev beam fixed

Den and I found the output beam of ETMY oplev was clipped the other day. Also I found the scattered beam of ITMY oplev was on the edge of the mirror inside the vacuum and it made more scattered lights.
ITMYoplev_before.jpg(before) -> IMG_0128.JPG(after)

  I fixed both of the clipped beam. But still the oplev feedback inject the noise into the arm. (red: oplev off, blue: oplev on)
   POYspe_oplev.pdf

  7807   Tue Dec 11 08:53:52 2012 SteveHowToPEMcables needs care

How NOT to:

The janitor can not clean in areas like this. He may only steps on these cables accidentally as he dust wiping our chambers.

Attachment 1: IMG_1839.JPG
IMG_1839.JPG
  7806   Mon Dec 10 22:34:34 2012 DenUpdateLSCbeam inside DRMI is clipping on PR3 Tip-Tilt

Quote:

 

That's OK, but its best to use standard notation. The power recycling gain is defined as the power incident on the BS divided by the power incident on the PRM from the laser side. You should also compare it with the PRC gain that you expect from mirror transmissions.

 I've made snapshots of PR2, PRM, ITMY and ITMX mirrors. Power buildup recycling gain (POWER BS / POWER PRM) was equal to 3-4.

PR2.bmp    PRM_LOCK.bmp    ITMY.png    ITMX.png

  7805   Mon Dec 10 16:28:13 2012 jamieOmnistructureComputersprogressive retrieval of online data now possible with the new NDS2 client

Leo fixed an issue with the new nds2-client packages that was preventing it from retrieving online data.  It's working now from matlab, python, and octave.

Here's an example of a dataviewer-like script in python:

#!/usr/bin/python

import sys
import nds2
from pylab import *

# channels are command line arguments
channels = sys.argv[1:]

conn = nds2.connection('fb', 8088)

fig = figure()
fig.show()
for bufs in conn.iterate(channels):
    fig.clf()
    for buf in bufs:
        plot(buf.data)
    draw()

  7804   Mon Dec 10 10:13:41 2012 DenUpdateLSCbeam inside DRMI is clipping on PR3 Tip-Tilt

 

That's OK, but its best to use standard notation. The power recycling gain is defined as the power incident on the BS divided by the power incident on the PRM from the laser side. You should also compare it with the PRC gain that you expect from mirror transmissions.

  7803   Mon Dec 10 03:02:03 2012 DenUpdateLSCbeam inside DRMI is clipping on PR3 Tip-Tilt

Quote:

 

 Some explanation of how you define power buildup please. Also some plots showing the evidence.

 I think about power buildup as a ratio of the power in the cavity when it is locked and unlocked = (POYDC_LOCKED - POYDC_OFFSET) / (POYDC_UNLOCKED - POYDC_OFFSET). I do not multiply this number by PRM transmission.

POYDC_OFFSET = -0.006

POYDC_UNLOCK = 0.063

For example, on the plot below power buildup is 15.

PRCL_LOCK.png

  7802   Sun Dec 9 17:51:53 2012 ranaUpdateLSCbeam inside DRMI is clipping on PR3 Tip-Tilt

 

 Some explanation of how you define power buildup please. Also some plots showing the evidence.

  7801   Sun Dec 9 01:21:51 2012 DenUpdateLSCbeam inside DRMI is clipping on PR3 Tip-Tilt

Quote:

 Today I wanted to check that AS and REFL beams are real and contain proper information about interferometer. For this I locked YARM using AS55_I and REFL11_I. Then I compared spectrum with POY11_I locking. Everything is the same. I've also adjusted phase rotations of AS55 (0.2 ->24) and REFL11 (-34.150 -> -43).

 I studied more carefully beam path inside DRMI using PRM face camera and found that beam is clipping on PR3 edge.

Step 1: PRCL LOCK, MICH LOCK, power build up 30.

Note: left is right and vice versa on the PRM camera

prcl_lock.mjpg

 Step 2: PRLC - UNLOCK, MICH - LOCK, PRM is still aligned. Right photo is AS port. I've slightly misaligned ITMs such that disturbance of AS beam is clearly seen.

PRM_UNLOCK.bmp       AS_UNLOCK.bmp

 

Step 3: PRCL - UNLOCK, MICH - LOCK, PRM misalined in yaw such such that the beam LASER -> PRM -> PR2 -> PR3 -> BS -> ITMX -> BS -> PR3 -> PR2 -> PRM -> PR2 -> PR3 is completely clipped on the TT edge. AS beam is now not clipped.

PRM_MISALIGN.bmp    AS_MISALIGN.bmp

So the conclusion is that when PRC is not locked and beam is thin, it can avoid clipping. When PRC locked, beam size grows and it starts to clip. I think we need to move the mount next to PR3 because of it we to not have enough space to align the TT.

Step 4: PSL shutter is closed.

PRM_BLOCK.bmp

  7800   Sat Dec 8 04:12:38 2012 DenUpdateLSCprcl

 Today I wanted to check that AS and REFL beams are real and contain proper information about interferometer. For this I locked YARM using AS55_I and REFL11_I. Then I compared spectrum with POY11_I locking. Everything is the same. I've also adjusted phase rotations of AS55 (0.2 ->24) and REFL11 (-34.150 -> -43).

Then I've locked MICH and aligned EMTs such that ASDC was close to zero. Then I locked PRCL and aligned PRM. Power buildup was 50. 

IMG_0118.JPG

  7799   Fri Dec 7 20:34:31 2012 AyakaUpdateSUSoplevs as acoustic and seismic noise couplers at high frequencies

I will just leave the picture of spectrum that shows the injected acoustic sound effects due to the oplevs.

 YARM_acous_oplev.pdf

red line: POY error without oplev feedback nor acoustic noise
blue line: POY error without oplev feedback but with acoustic noise
brown line: POY error with oplev feedback but without acoustic noise
green line: POY error with oplev and acoustic noise

You can see there is noise only at green line around 70 - 100 Hz. And it does not look like the acoustic signal is injected directly to the arms but the acoustic sound couples to the original noise source.

  7798   Fri Dec 7 19:24:43 2012 Den, AyakaUpdateSUSoplevs as acoustic and seismic noise couplers at high frequencies

We've provided acoustic excitation using speakers on the AS table and saw that PSD of YARM feedback signal increased in the frequency range 50 - 100 Hz. Meanwhile, XARM feedback signal did not change. Moreover, YARM noise is much higher at these frequencies compared to XARM.

The problem was with YARM oplev servos. Both ITMY and ETMY produced noise to YARM length. ITMY oplev signal had a huge resonance at 55 Hz. We measured coherence with accelerometers, it was 0.8. It turned out that one of the mirror mounts was not fixed in the oplev path. When we fixed it, noise has gone.

Note: speakers were on AS table but mirror mounts could steel feel it on ITMY table.

IMG_0111.JPG         YARM.png

Then we had a look on ETMY table. We saw a mirror on suspiciously long mirror mount that was used in the ETMY oplev path. We slightly kicked long mount with a small screwdriver and YARM control signal went up with resonance at 100 Hz.

IMG_0112.JPG        ETMY_olreso_copy.png

  7797   Fri Dec 7 02:02:23 2012 DenUpdateAlignmentc1ass

I suppose that there is also a similar system possible to align the arms in a continuous way; i.e. low level drives and very low bandwidth. Also something fast / slow for the the DRMI.

 c1ass was really useful today when we slowly aligned PZT and servo kept arms aligned to the input beam. I think it is possible to automate phase and matrix measurements. DRMI servo will be very useful.

Today I tried to investigate the mode in PRCL and MICH. I locked them but power build-up was only 27. The beam on the POP camera looked like interference of 00 mode and a long strip of fringes. (I wanted to make videocaptures but script is not working - the problem is that it is looking for /usr/lib/*.so.4 libraries but they were updated to *.so.5, I made a few links .so.4 -> .so.5 but this kept going for many libraries, so this should be fixed in a better way). 

We looked at PRM and BS faces and they had the same shape - interference of a circle with a strip. There were also a lot of bright spots all over the frames. Loops were closed and circle was not moving. Strip was oscillating at ~1Hz and also its position significantly changed with alignment. Looking at PRM face camera we made a conclusion that the length of the strip is ~5 cm and width ~1cm. Interesting that strip has plenty of power - approximately 10 times of transmitted beam when cavity is not locked. As a result POYDC was oscillating at the same frequency as a strip.

  7796   Fri Dec 7 00:08:39 2012 ranaUpdateAlignmentc1ass

 

 This looks like a good performance tuning for these. It would be good if you can codify this procedure in the wiki so that even unexperienced people can tune up the system after reboots or vacuum work.

Is it possible to have some python scripts automatically measure and set the phases and matrices? If so, can we also run them iteratively so that after the second run we can confirm that they have converged? Then the script can output a short report of numbers telling us how well the system is now tuned.

I suppose that there is also a similar system possible to align the arms in a continuous way; i.e. low level drives and very low bandwidth. Also something fast / slow for the the DRMI.

  7795   Thu Dec 6 02:55:46 2012 DenUpdateAlignmentc1ass

Today I've set c1ass model to improve alignment of X and Y arms. I've added all measured parameters to ASS scripts. I've also added a script to c1ass.adl that downloads calculated OFFSETs to corresponding ASC filter banks and blocks outputs. It should be called after alignment convergence.

XARM phase rotation and sensing matrix

Demodulator Phase rotation, degrees
ETM_P_T -5
ETM_Y_T -10
ITM_P_T -62
ITM_Y_T 163

                                                           

I-quadrature ETM_PIT ETM_YAW ITM_PIT ITM_YAW
ETM_P_T

1.0000  

-0.6116 -0.560 0.5660
ETM_Y_T -0.1600 1.0000 0.2310 -0.3979
ITM_P_T -0.0794 0.4960 1.0000 -0.8791
ITM_Y_T 0.0624 -0.3903 -0.4787 1.0000

 

Output gains were (-0.5, 0.5, -0.25, -0.25). XARM Gain was set to 0.5.

xarm_ass.png

YARM phase rotation and sensing matrix

Demodulator Phase rotation, degrees
ETM_P_T 10
ETM_Y_T 0
ITM_P_T 107
ITM_Y_T -35

 

I-quadrature ETM_PIT ETM_YAW ITM_PIT ITM_YAW
ETM_P_T

1.0000  

0.3899 -0.515  -0.0309
ETM_Y_T 0.1017 1.0000 -0.3143 -0.5269
ITM_P_T -0.3505 0.0565 1.0000 -0.0945
ITM_Y_T -0.2085 -0.3607 0.6042 1.0000

Output gains were (-0.25, 0.25, 0.7, -0.7). YARM Gain was set to 0.8.

yarm_ass.png

  7794   Wed Dec 5 17:38:41 2012 RijuHowTo Photodiode transimpedance

I have started making the circuit to measure the transimpedance for the photodiode PDA10CF using Jenne's laser. I will continue it tomorrow.

  7793   Wed Dec 5 16:54:29 2012 jamieOmnistructureComputersnew (beta) version of NDS2 installed on control room machines

Quote:

NDS2 is not designed for non DQ channels - it gets data from the frames, not through NDS1.

For getting the non-DQ stuff, I would just continue using our NDS1 compatible NDS mex files (this is what is used in mDV).

The NDS2 protocol is not for non-DQ, but the NDS2 client is capable of talking both the NDS1 and NDS2 protocols.

fb:8088 is an NDS1 server, so the client is talking NDS1 to fb.  It should therefore be capable of getting online data.

It doesn't seem to be seeing the online channels, though, so I'll work with Leo to figure out what's going on there.

The old mex functions, which like I said are now available, aren't capable of getting online data.

  7792   Wed Dec 5 09:53:01 2012 ranaUpdateWienerFilteringThe microphones and the speaker on the AP table

  Don't try to re-invent the mic mount: just copy the LIGO mic mount for the first version.

  7791   Wed Dec 5 09:42:46 2012 ranaOmnistructureComputersnew (beta) version of NDS2 installed on control room machines

NDS2 is not designed for non DQ channels - it gets data from the frames, not through NDS1.

For getting the non-DQ stuff, I would just continue using our NDS1 compatible NDS mex files (this is what is used in mDV).

  7790   Wed Dec 5 03:25:32 2012 DenUpdateASCdithering

I wanted to center beams on the XARM cavity mirrors using c1ass model. I've run XARM setup script and then turned dithering on. Cavity went out of lock because calculated offsets were incorrect.

I was using TRX only and calculated rotation phases for ITM and ETM pitch and yaw. For this I've added a low pass filter into Q-quadrature bank and made DC value at the output to be zero by adjusting the phase. I've put gains (+1 or -1)  in the I quadrature such that output was positive.

Then I've set the sensing matrix to identity as I decided to deal with separate loops. Of coarse, they are mixed by the cavity, but at least in the control system they are distinguished. Old matrix summed error signals in one degree of freedom from both mirrors. This makes more sense but still not precise because coils are not ideally diagonalized.

Then I've adjusted gains for control loop for every degree of freedom. I've ended up with (0.1; 0.1; 0.1; -0.1). I did not use large gains as I wanted slow convergence because of the demodulation low-pass filter time response constant of 20 sec. Coupling (I quadrature) was reduced from (0.9, 0.3, 2.4, 1.2) to zeros (0-0.1) in ~5 minutes, TRX increased from 0.73 to 0.90.

There is one thing that I do not understand yet. I think controllers should minimize angle -> length coupling that is proportional to I-quadrature if phase is correct. But phase depends on alignment and when the feedback loops are on, phase drifts. I could see it during my measurement. But I did not find any script that smoothly tunes phase such that coupling is all in I-quadrature. I guess this is not hard to set a gradient descent algorithm that minimizes DC value of Q-quadrature. Or how this is usually done?

  7789   Wed Dec 5 01:06:22 2012 AyakaUpdateWienerFilteringThe microphones and the speaker on the AP table

 In order to see the acoustic coupling on arm signals, I set 6 microphones and the speaker on the AP table. The microphones are not seismically isolated for now.
I have a signal generator under the AP table.

DSC_4956.JPGDSC_4961.JPG

When I played the 43 Hz triangular wave sound, I could see some coherence between POY error signal and microphones even though there is no peak in POY.

APsig.pdf

To Do:

  • Try to subtract the acoustic signal and see with which microphone the acoustic signal can be subtracted best. But how can I find whether the signal is subtracted or not? Is coherence information enough?
  • Make circuits for microphones to come to 40m.
  • Make suspension systems for microphones. One idea is that the microphones should be suspended from bridges which is to be put around at the top of the tables since there is no space for stacks for each microphones.
  • Prepare a new ADC.
  • Perform the same measurements at the other tables, such as POX and POY.
  7788   Tue Dec 4 23:08:46 2012 DenOmnistructureComputersnew (beta) version of nds2 installed on control room machines

Quote:

I've installed the new nds2 packages on the control room machines.

 I've tried new nds2 Java interface in Matlab. Using findChannels method of the connection class I see only slow, DQ and trend channels. I could even download data online using iterate method. When it will be possible to do the same with fast non-DQ channels?

>> conn = nds2.connection('fb', 8088);
>> conn.iterate({'C1:LSC-XARM_OUT'})
??? Java exception occurred:
java.lang.RuntimeException: No such channel.
    at nds2.nds2JNI.connection_iterate__SWIG_0(Native Method)
    at nds2.connection.iterate(connection.java:91)

  7787   Tue Dec 4 21:57:04 2012 DenUpdateSUSTTs are ready

Quote:

Please leave here what was the instruction by Bran and Suresh so that the other people can redo it sometime later!

 The connectors can be plugged into the BOSEMs if we loosen the two screws which hold down the mini-D connector and the flex circuit.  Tighten the screws after the connector is pluged in.

  7786   Tue Dec 4 20:38:51 2012 jamieOmnistructureComputersnew (beta) version of nds2 installed on control room machines

I've installed the new nds2 packages on the control room machines.

These new packages include some new and improved interfaces for python, matlab, and octave that were not previously available. See the documentation in:

  /usr/share/doc/nds2-client-doc/html/index.html

for details on how to use them.  They all work something like:

  conn = nds2.connection('fb', 8088)
  chans = conn.findChannels()
  buffers = conn.fetch(t1, t2, {c1,...})
  data = buffers(1).getData()

NOTE: the new interface for python is distinct from the one provided by pynds.  The old pynds interface should continue to work, though.

To use the new matlab interface, you have to first issue the following command:

   javaaddpath('/usr/lib/java')

I'll try to figure out a way to have that included automatically.

The old Matlab mex functions (NDS*_GetData, NDS*_GetChannel, etc.) are now provided by a new and improved package.  Those should now work "out of the box".

  7785   Tue Dec 4 20:13:55 2012 KojiUpdateSUSTTs are ready

Please leave here what was the instruction by Bran and Suresh so that the other people can redo it sometime later!

  7784   Tue Dec 4 18:53:50 2012 AyakaUpdateAlignmentBS chamber

BS chamber seemed to be kicked again around 10:00 am today.

I moved PZT mainly in YAW and locked both arms. I adjusted the beam to be almost on the center of both ETM by sights.

BSchamb1204.pdf

  7783   Tue Dec 4 18:06:35 2012 DenUpdateSUSTTs are ready

 Using instructions from Bram and Suresh, I was able to plug in connectors to BOSEMs. Today I've tested electronics, everything works good. Jamie made an medm screen and channels for TTs. Sliders for pitch and yaw go from -100 to 100 counts. Calibration to angle is 1e-5 rad / count.

TTs are in the clean room waiting for installation.

IMG_0105.JPG    IMG_0108.JPG

  7782   Tue Dec 4 11:30:36 2012 ZachUpdatePSLPMC calibration for MC_F calibration

In order to have less unknown, you can calibrate the PMC PZT separately. Lock the PMC and take a transfer function from either the NPRO PZT input or the FSS AOM VCO input to the PMC control signal. The VCO is better, since the calibration should be much better known, but I am not sure what the current setup of the 40m PSL is, so I don't know if the FSS is normally locked.

Since you know the NPRO PZT or VCO actuation coefficients, you can assume the PMC loop (where the OL gain is high enough) is correcting for the frequency fluctuations. So, simply multiply the known coefficient by the transfer function to get the PMC PZT gain.

Then, you can re-do your PMC PZT sweep measurement and be confident of the calibration. The FSR must be right, so you can get the finesse with confidence.

Quote:

Hmm... Does anyone find falses in my measurement?
If not, the finesse can be 4 times smaller than the value which was 5 years ago?

 

  7781   Tue Dec 4 10:51:10 2012 SteveUpdateSUSPRM damping restored

PRM oplev gains set to zero from PIT 0.15 and YAW -0.3 and damping restored

  7780   Tue Dec 4 10:44:36 2012 SteveUpdatePEMshort power outage

 

CALIFORNIA INSTITUTE OF TECHNOLOGY

                 FACILITIES MANAGEMENT

 

**PLEASE POST**

 

 

Building:         Campus

 

Date:             Thursday October 04,2012

 

This morning at 2:17 a.m. much of the City of Pasadena including our Campus experienced a electric power sag of short duration, approximately 1/10 of a second. The cause was a fault on one of Pasadena’s 17KV circuits. Some sensitive equipment have been impacted.

                 

Contact:          Mike Anchondo x-4999

  7779   Tue Dec 4 01:43:37 2012 ranaUpdatePSLPMC calibration for MC_F calibration

  If you can't find the PMC sidebands in the transmission, its because the SNR is too small.

It may be a better idea to look at the PMC error signal, since the DC signal there is suppressed by the demodulation.

  7778   Mon Dec 3 17:04:12 2012 AyakaUpdatePSLPMC calibration for MC_F calibration

In order to calibrate MC_F signal, I need to know the calibration value from thorlab's PZT driver to laser frequency.
The calibration value should be ~ 15MHz/V (the PZT driver has 15 gain, and the laser has the calibration value of ~ 1MHz/V according to the laser spec sheet), but I want to confirm it.
This can be measured by sweeping the input voltage of the PZT driver and see the transmission signal from unlocked PMC.
 

1. Response of PMC transmission when the signal is inputted to laser PZT

I inputted 0.2 Hz triangular wave with 5Vamp and 2.5V offset into the PZT driver and see the transmission signal from PMC. After the PZT driver and before the laser, there is an analog low pass filter but its cut off frequency is 1 Hz so I did not take it into consideration. 
DSC_4955.JPGsweep_PZTdriver.png(TEK00000.CSV, TEK00001.CSV in the zip file)

I could not the side-band resonances. I guess it was because the generated signal is not big enough (but still the maximum range of the signal generator.)
Therefore, in order to calibrate the input voltage to the frequency, I need to know finesse or FWHM frequency.
 

2. Responce of PMC transmission when the voltage of PZT on the PMC is swept

In order to measure the finesse and FWHM frequency, I also swept the PMC PZT voltage with the DC offset slider at the FSS.adl and tried to measure the finesse of PMC. (reference: elog #904)
PMC_PZT_FSR.pdftrans_fit.png(PMC-PZTcal_121203.xml in the zip file)

The result of fitting:

V_FSR (the PZT voltage difference between the 2 resonances) ~ 63 +/- 7 V (= 731MHz (given))
V_FWHM (the PZT voltage to sweep FWHM) ~  0.32 +/- 0.04 V (~ 3.7 MHz)
Finesse ~ 200 +/- 30

However, this finesse value is much smaller than the value on the Wiki, 800. (Manasa showed me.)
V_FSR is comparable to the result Rana got at the referenced elog. But I am not sure about the V_FWHM because it is hard to figure out how large the PZT voltage changed from the template file (PMC-PZTcal.xml).
Are those mode wrong? But if so, where is the correct mode resonances? I think they should be visible...
 

3. Calibration value 

When I know the FWHM frequency, I can calibrate the input on the PZT driver into laser frequency.

The results are:

if I take the finesse of 800 and FSR of 731 MHz (the values on the Wiki): ~5.0 MHz/V
if I take the finesse of 200 and FSR of 731 MHz (the measured value): ~20.0 MHz/V

Actually, the measured value is closer to the value calculated from the spec sheet.

Hmm... Does anyone find falses in my measurement?
If not, the finesse can be 4 times smaller than the value which was 5 years ago?

Attachment 5: PMC-PZTcal_121203.zip
  7777   Mon Dec 3 16:37:09 2012 SteveUpdatePEMEM 172 microphones ordered

Quote:

I've put EM 172 microphones inside Steve's isolation box to measure their noise. I've attached mics to each other and aligned them using the tape.

At low frequencies (below 1 Hz) the noise is limited by ADC as there is a 10 Hz high-pass filter inside mic readout box.

ADC noise is measured by splitting the signal from 1 mic into 2 ADC channels.

em172.png

 BT EM172 microphones  are ordered.

ELOG V3.1.3-